OpenStudy (anonymous):

This is a multiple step exponential growth problem, dealing with zombies, help would be amazing!! Zombies have infected WSU-V after 5 zombies escaped confinement in the science lab at 3:14 AM six hours after their escape there were 72 zombies on campus. Assuming an exponential increase write a function p(h) that represents the population (p) in terms of time in hours (h) If some one would be willing to walk me through this... this is just the first part

OpenStudy (anonymous):

what method would you like to use? we can use \(e\) or do it an easier way

OpenStudy (anonymous):

are you supposed to write something like \[p(h)=P_0e^{kh}\] ?

OpenStudy (anonymous):

it needs to be in terms of an exponential growth function, I'm assuming something along the lines of p(h)=(1+.05)^x? however if we could find an equation that way, it would be just fine

OpenStudy (anonymous):

sorry, that would be P=C(1+r)^t I have no idea where i got that other equation from...

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

C being the initial value, r being the rate, and t being time?

OpenStudy (anonymous):

so we are not using \(e\) right?

OpenStudy (anonymous):

I would presume not, however I am open to ideas seeing as i have attempted this multiple times..

OpenStudy (anonymous):

\(C\) is given to you, it is what you start with, namely \(5\)

OpenStudy (anonymous):

now \(1+r\) is also easy enough, \(r\) is the increase as a decimal (not a percent) easiest way to find \(1+r\) is to take \(\frac{72}{5}\) or \(14.4\)

OpenStudy (anonymous):

but this takes \(6\) hours, not 1 hour, so this method would give you \[\large P(h)=5\times (14.4)^{\frac{h}{6}}\]

OpenStudy (anonymous):

what is the h/6?

OpenStudy (anonymous):

if you want just an \(h\) in the exponent, rather than a \(\frac{h}{6}\) use for your base \((14.4)^{\frac{1}{6}}\) instead of \(14.4\) you will need a calculator to find it

OpenStudy (anonymous):

it is \(\frac{h}{6}\) because it takes 6 hours to increase by a factor of \(14.4\) not one hour

OpenStudy (anonymous):

how did you come to h/6 though? what is h? and why does h/6=1?

OpenStudy (anonymous):

they asked you to use \(h\) as the variable in the exponent, i.e. the time measured in "hours" if you want to use \(t \)for time, that is fine too, it is the same thing

OpenStudy (anonymous):

oh is it because it took 6 hours for them to reach the current population so that is your reference equation? Then you substitute 6 on top for h or hours?

OpenStudy (anonymous):

okay, I think i understand that, so if it were to be 9 hours instead of 6, it would be h/9?

OpenStudy (anonymous):

right i used \(h\) because it said to write \(p(h)\) yes, exactly!

OpenStudy (anonymous):

but if you just want the exponent to be \(h\) rather than \(\frac{h}{6}\) compute \[14.4^{\frac{1}{6}}\] and use that as your base then the exponent will just be \(h\) lets find it

OpenStudy (anonymous):

looks like it is about \(1.56\) http://www.wolframalpha.com/input/?i=%2872%2F5%29^%281%2F6%29

OpenStudy (anonymous):

so an alternate answer would be \[p(h)=5\times (1.56)^h\]

OpenStudy (anonymous):

great, thanks! this question also asks me how many zombies will there be at 4:05 PM? and At what time will there be 100,000 zombies? if it is the next day, please indicate it. what values would you have to substitute in? (that's my own question)

OpenStudy (anonymous):

also how did you arrive at 1.56? and why would you use h instead of h/6??

OpenStudy (anonymous):

by the law of exponents \[\large 14.4^{\frac{h}{6}}=(14.4^{\frac{1}{6}})^h\]

OpenStudy (anonymous):

so if you want just an \(h\) up in the sky, compute \((14.4)^{\frac{1}{6}}\) and use that as your base

OpenStudy (anonymous):

i used a calculator to get it it really makes no difference, just if for some reason they want the exponent to be \(h\) instead of \(\frac{h}{6}\) that is what you have to do

OpenStudy (anonymous):

okay, I think I get that, but how would you find the number of zombies at 4:05 PM? you would have to find the time difference and substitute that for t?

OpenStudy (anonymous):

yes, but you would also have to convert the minutes to hours, since your units are hours kind of a pain, but that is what you need for this

OpenStudy (anonymous):

so a 12 hour and 50 minute difference would be a 12 5/6 exponent?

OpenStudy (anonymous):

oh right am to pm , didn't catch that good call

OpenStudy (anonymous):

Thanks, almost missed it myself :) so your equation would be: 5(14.4)^(12 5/6)/6 or just 5(14.4)^12 5/6?

OpenStudy (anonymous):

guess you are rounding, that is fine

OpenStudy (anonymous):

first one if you are using the base as 14.4

OpenStudy (anonymous):

i would convert the mixed number to a fraction first

OpenStudy (anonymous):

77/5

OpenStudy (anonymous):

actually 77/6

OpenStudy (anonymous):

then divide by 6 to get 77/36

OpenStudy (anonymous):

right, and i was rounding so that would actually be incorrect, my bad!!

OpenStudy (anonymous):

then compute \[5\times (14.4)^{\frac{77}{36}}\] i don't thing the rounding error is so bad

OpenStudy (anonymous):

i get about 1500

OpenStudy (anonymous):

why is it 77 over 30?

OpenStudy (anonymous):

oh, never mind..

OpenStudy (anonymous):

\(36\) not \(30\) that is because we are using \[\large p(h)=5\times (14.4)^{\frac{h}{6}}\] with \(h=\frac{77}{6}\)

OpenStudy (anonymous):

1513 zombies, if we round up, although i suppose you could have half a zombie...

OpenStudy (anonymous):

close enough for government work not sure how picky your teacher or whoever is

OpenStudy (anonymous):

Oh he's very picky, CDO as he likes to say, its OCD with the letters in alphabetical order.

OpenStudy (anonymous):

so for the second part (how long until 100,000) you would substitute in values until you get to 100,000?

OpenStudy (anonymous):

well then i guess you shouldn't round, even with the minutes for that you either need to grind it til you find it or use logarithms depending on if you got to that or not, i am going to guess not

OpenStudy (anonymous):

Well i simply added on a minute so 50 became 51 and that answer should be correct down to the decimal 1512.851644 zombies

OpenStudy (anonymous):

k leave it then looks good

OpenStudy (anonymous):

although this is actually a test i am re-doing I'll probably end up getting completely different problems tomorrow knowing him.. was i correct in the process for the second part though?

OpenStudy (anonymous):

yes, just keep checking this was on a test?!

OpenStudy (anonymous):

Yes, it was, good or bad?

OpenStudy (anonymous):

hard i would say, and annoying with hours and minutes etc

OpenStudy (anonymous):

It was not fun. At all. and he ruined zombies for me, but thank you so much, I really appreciate your help!

OpenStudy (anonymous):

yw